14. Find the equation of the tangent line to the graph f(x) = e24 ln(22) at the point (1,0).
Answer #1

f(x) = 82 m (22) . (60) - 2e2ma f(a)= 2227 + 2m2. 202% at (LO), fW-.2.? ! *2ml, 2e! Co2e40 - 282 ea of tangent y-o=202(x-1)